Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Siblings

Expert replies
by neerajkumar1_1 » Mon Sep 13, 2010 3:53 am
In a room filled with 7 people, 4 people have exactly 1 sibling in the room and 3 people have exactly 2 siblings in the room. If two individuals are selected from the room at random, what is the probability that those two individuals are NOT siblings?
5/21
3/7
4/7
5/7
16/21

This question has been answered before on the forum...
but I cant understand the concept of siblings here...

4 people have exactly 1 sibling means what... and 3 people have exactly 2 siblings means what...

Plz explain... thanks...
Join the discussion
Source: — Problem Solving |

by Jim@Grockit » Mon Sep 13, 2010 5:16 am
neerajkumar1_1 wrote:In a room filled with 7 people, 4 people have exactly 1 sibling in the room and 3 people have exactly 2 siblings in the room. If two individuals are selected from the room at random, what is the probability that those two individuals are NOT siblings?
5/21
3/7
4/7
5/7
16/21

This question has been answered before on the forum...
but I cant understand the concept of siblings here...

4 people have exactly 1 sibling means what... and 3 people have exactly 2 siblings means what...

Plz explain... thanks...
"Siblings" means "brothers or sisters;" for simplicity's sake, consider replacing "sibling" in your head with "relative." It means that 4 people have one relative in the room, and 3 people have two relatives in the room. Does that help?
Join the discussion

by neerajkumar1_1 » Mon Sep 13, 2010 5:23 am
Jim@Grockit wrote:
neerajkumar1_1 wrote:In a room filled with 7 people, 4 people have exactly 1 sibling in the room and 3 people have exactly 2 siblings in the room. If two individuals are selected from the room at random, what is the probability that those two individuals are NOT siblings?
5/21
3/7
4/7
5/7
16/21

This question has been answered before on the forum...
but I cant understand the concept of siblings here...

4 people have exactly 1 sibling means what... and 3 people have exactly 2 siblings means what...

Plz explain... thanks...
"Siblings" means "brothers or sisters;" for simplicity's sake, consider replacing "sibling" in your head with "relative." It means that 4 people have one relative in the room, and 3 people have two relatives in the room. Does that help?
Well thats not the issue..
I saw the solns...
How does one sibling for 4 people relate to 2 pairs and so on for 3 people with 2 siblings...
I am just getting thoroughly confused and am not able to visualize the problem...

Hope u can provide me with a little more detail...
Join the discussion

by neerajkumar1_1 » Tue Sep 14, 2010 2:54 am
hi guyz... can someone plz provide me with a solution and including how we make the sets of siblings...
Join the discussion

by kmittal82 » Tue Sep 14, 2010 3:10 am
>people have exactly 1 sibling

Let them be A1 A2 B1 B2

Siblings: A1-A2 and B1-B2

>3 people have exactly 2 siblings in the room

C1 C2 C3

Siblings: C1-C2-C3

To further simplify, you can say the people are A A B B C C C, but keep in mind when doing any selections, the order will matter (so use a permutation)
Join the discussion

by neerajkumar1_1 » Tue Sep 14, 2010 3:14 am
kmittal82 wrote:>people have exactly 1 sibling

Let them be A1 A2 B1 B2

Siblings: A1-A2 and B1-B2

>3 people have exactly 2 siblings in the room

C1 C2 C3

Siblings: C1-C2-C3

To further simplify, you can say the people are A A B B C C C, but keep in mind when doing any selections, the order will matter (so use a permutation)
hi,
In the fist set which u made..
where it is given that 4 people have one sibling...

how do u make 2 relations
Siblings: A1-A2 and B1-B2

I am just failing to interpret the question.. :(
Join the discussion

by GMATGuruNY » Tue Sep 14, 2010 3:51 am
neerajkumar1_1 wrote:In a room filled with 7 people, 4 people have exactly 1 sibling in the room and 3 people have exactly 2 siblings in the room. If two individuals are selected from the room at random, what is the probability that those two individuals are NOT siblings?
5/21
3/7
4/7
5/7
16/21

This question has been answered before on the forum...
but I cant understand the concept of siblings here...

4 people have exactly 1 sibling means what... and 3 people have exactly 2 siblings means what...

Plz explain... thanks...
Let's say that the 7 people are ABCDEFG.

4 people have exactly 1 sibling:
Let's say that A and B are siblings and that C and D are siblings.
This means:
A has 1 sibling (B).
B has 1 sibling (A).
C has 1 sibling (D).
D has 1 sibling (C).

3 people have exactly 2 siblings:
Let's say that E, F and G are all siblings of each other.
This means:
E has 2 siblings (F and G).
F has 2 siblings (E and G).
G has 2 siblings (E and F).

Total number of sibling pairs = 5: AB, CD, EF, EG, FG.
Total number of pairs that can be formed from 7 people: 7C2 = 21.
P(sibling pair) = 5/21
P(not sibling pair) = 1 - 5/21 = 16/21.

Does this help?
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

by neerajkumar1_1 » Tue Sep 14, 2010 4:09 am
GMATGuruNY wrote:
neerajkumar1_1 wrote:In a room filled with 7 people, 4 people have exactly 1 sibling in the room and 3 people have exactly 2 siblings in the room. If two individuals are selected from the room at random, what is the probability that those two individuals are NOT siblings?
5/21
3/7
4/7
5/7
16/21

This question has been answered before on the forum...
but I cant understand the concept of siblings here...

4 people have exactly 1 sibling means what... and 3 people have exactly 2 siblings means what...

Plz explain... thanks...
Let's say that the 7 people are ABCDEFG.

4 people have exactly 1 sibling:
Let's say that A and B are siblings and that C and D are siblings.
This means:
A has 1 sibling (B).
B has 1 sibling (A).
C has 1 sibling (D).
D has 1 sibling (C).

3 people have exactly 2 siblings:
Let's say that E, F and G are all siblings of each other.
This means:
E has 2 siblings (F and G).
F has 2 siblings (E and G).
G has 2 siblings (E and F).

Total number of sibling pairs = 5: AB, CD, EF, EG, FG.
Total number of pairs that can be formed from 7 people: 7C2 = 21.
P(sibling pair) = 5/21
P(not sibling pair) = 1 - 5/21 = 16/21.

Does this help?
it surely does... phew!!!!>.. I was like so frustrated...
thanks a lot...
Join the discussion